40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16997     Entry time: Wed Jul 13 12:49:25 2022
Author: Paco 
Type: Summary 
Category: SUS 
Subject: SUS frozen 

[Paco, JC, Yuta]

This morning, while investigating the source of a burning smell, we turned off the c1SUS 1X4 power strip powering the sorensens. After this, we noticed the MC1 refl was not on the camera, and in general other vertex SUS were misaligned even though JC had aligned the IFO in the morning to almost optimum arm cavity flashing. After a c1susaux modbusIOC service restart and burt restore, the problem persisted.

We started to debug the sus rack chain for PRM since the oplev beam was still near its alignment so we could use it as a sensor. The first weird thing we noticed was that no matter how much we "kicked" PRM, we wouldn't see any motion on the oplev. We repeatedly kicked UL coil and looked at the coil driver inputs and outputs, and also verified the eurocard had DC power on which it did. Somehow disconnecting the acromag inputs didn't affect the medm screen values, so that made us suspicious that something was weird with these ADCs.


Because all the slow channels were in a frozen state, we tried restarting c1susaux and the acromag chassis and this fixed the issue.

ELOG V3.1.3-